OpenStudy (anonymous):

Ok I honestly have no clue how to do this please can someone explain to me what I have to do? Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y-1=4(x-3) y+1=4(x+3) y-3=4(x+1) y+3=4(x-1)

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] (x1,y1) m=slope plug in the info

OpenStudy (anonymous):

y-3=4(x+1)?

pooja195 (pooja195):

you have your X and Y values mixed up \(\huge\color{red}{(x_1,y_1)}\)

OpenStudy (anonymous):

I have -3 as x and -1 as y? ohh wait I see what you mean now it should be y-1=4(x-3)?

OpenStudy (anonymous):

would the 1 also become positive ?

pooja195 (pooja195):

\(\huge~\color{red}{y+1=4(x+3) }\)

OpenStudy (anonymous):

oh ok i get it Thank you sooo much!

OpenStudy (anonymous):

ok this is another one I have no clue how to do .....well I think I just dont know what format to use If f(x) = 2x2 - 30, find f(4).

pooja195 (pooja195):

repost that :) also you are welcome! \(\huge\color{orange}{Welcome~To~OpenStudy!!~:)}\)

OpenStudy (anonymous):

ok this is another one I have no clue how to do .....well I think I just dont know what format to use If f(x) = 2x2 - 30, find f(4).

pooja195 (pooja195):

On a difrent post meaning close this and post again

OpenStudy (anonymous):

ohh ok